OpenStudy (anonymous):

how to graph and solve 10x-3y=15

OpenStudy (asevilla5):

okay so first you have to solve the inequality. 10x-3y=15 you want the y=mx+b format so.. you subtract -10x on both sides to get -3y=-10x+15 then later on your divide -3 on both sides to later get y=10/3+3 the 10/3 is simplified. to graph this